OpenStudy (noodles2061):

a triangle has side lengths of 24 in, 32, in, and 40 in. is the triangle acute, obtuse, or right?

OpenStudy (raffle_snaffle):

use pythag a^2 + b^2 = c^2

OpenStudy (noodles2061):

but how do i use that to see what kind of angle it is?

OpenStudy (ask_if_im_a_ghost):

If c^2<a^2+b^2 the the triangle is acute c^2 = a^2+b^2 it's right c^2 > a^2+b^2 obtuse Where a, b, c are the different sides of the triangle

OpenStudy (welshfella):

a^2 > b^2 + c^2 where a is longest side - if that is true then the triangle is obtuse

OpenStudy (ask_if_im_a_ghost):

In my equations, c is the longest side

OpenStudy (welshfella):

- as Ghost siad

OpenStudy (noodles2061):

okay, thank you!!!

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\)\(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)